OpenStudy (anonymous):

What is the distance points between the points (-1,-3) and (-2,-2)

OpenStudy (solomonzelman):

\(\large\color{blue}{ \rm {distance~~between~~(x_1,y_1)~~and~~(x_1,y_2)}=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2} }\)

OpenStudy (solomonzelman):

And similarly here \(\normalsize\color{blue}{ \rm {distance~between~(-1,-3)~and~(-2,-2)}=\tt{ \sqrt{(-1--2)^2+(-3--2)^2} }}\)

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ =\tt{ \sqrt{(-1+2)^2+(-3+2)^2} }}\) \(\normalsize\color{blue}{ =\tt{ \sqrt{1^2+(-1)^2} }}\) \(\normalsize\color{blue}{ =\tt{ \sqrt{1+1} }}\) \(\normalsize\color{blue}{ =\tt{ \sqrt{2} }}\) \(\normalsize\color{blue}{ (~\tt{ \sqrt{2}~≈ ~1.414) }}\)

OpenStudy (anonymous):

Wait then what's the answer

OpenStudy (solomonzelman):

You can write √2, or approximate √2, to 1.141

OpenStudy (solomonzelman):

I''ll post some instructions, do you mind ?

OpenStudy (anonymous):

No I don't mind

OpenStudy (anonymous):

Answers A: √26 B:√13 C: 4√5 D:6.5 Which one is it

OpenStudy (solomonzelman):

\(\large\color{ blue }{\large {\bbox[5pt, cyan ,border:2px solid white ]{ \rm I~ }}}\)\(\large\color{ cyan }{\large {\bbox[5pt, green ,border:2px solid white ]{ \rm N }}}\)\(\large\color{ green }{\large {\bbox[5pt, red ,border:2px solid white ]{ \rm S }}}\)\(\large\color{ red }{\large {\bbox[5pt, purple ,border:2px solid white ]{ \rm T}}}\)\(\large\color{ purple }{\large {\bbox[5pt, yellow ,border:2px solid white ]{ \rm R}}}\)\(\large\color{ yellow }{\large {\bbox[5pt, magenta ,border:2px solid white ]{ \rm U}}}\)\(\large\color{ magenta }{\large {\bbox[5pt, teal ,border:2px solid white ]{ \rm C}}}\)\(\large\color{ teal }{\large {\bbox[5pt, brown ,border:2px solid white ]{ \rm T}}}\)\(\large\color{ brown}{\large {\bbox[5pt, goldenrod ,border:2px solid white ]{ \rm I~}}}\)\(\large\color{ goldenrod}{\large {\bbox[5pt, midnightblue ,border:2px solid white ]{ \rm O}}}\)\(\large\color{ midnightblue}{\large {\bbox[5pt, lightgreen ,border:2px solid white ]{ \rm N}}}\)\(\large\color{ lightgreen}{\large {\bbox[5pt, blue ,border:2px solid white ]{ \rm S}}}\) \(\normalsize\color{blue}{\rm {you~~~can~~~find~~~any~~~section~~~by:} }\) \(\normalsize \color{black}{\rm { 1) } }\) clicking "more subjects" (the top left corner, towards the center) \(\normalsize \color{black}{\rm { 2) } }\) typing in the name of the section (name of a section can be "history" Mathematics and other section for each unique area of studying). \(\normalsize\color{blue}{\rm {You~~~can~~~FAN~~~someone~~~by:} }\) \(\normalsize\color{black}{\rm { 1) } }\) clicking on the user's name \(\normalsize\color{black}{\rm { 2) } }\) clicking "fan" on the bottom right corner of user's box. \(\normalsize\color{black}{\rm { (~ } }\)You can also write a testimonial by clicking "add testimonial" and typing something in testimonial box.\(\normalsize\color{black}{\rm { ~) } }\) \(\normalsize\color{blue}{\rm {You~~~can~~~mention~~~someone~~~in~~~your~~~question} }\) (or in any other question) by writing @ in front of their name, just like I would mention you... LOOK ! .... @CALIFORNAx3 (they would get a notification that a user mentioned them in a reply) \(\normalsize\color{blue}{\rm {You~~~can~~~ message~~~someone~~~by:} }\) \(\normalsize\color{black}{\rm { 1) } }\) clicking on the someone's username \(\normalsize\color{black}{\rm { 2) } }\) clicking "Send Message" \(\normalsize\color{black}{\rm { 3) } }\) type your message press submit, and wait till the messages submits. \(\normalsize\color{blue}{\rm {You~~~can~~~ see~~~if~~~a~~~user~~~is~~~offline~~~by} }\) Clicking on the user's username, and below his picture you will see (in light blue) “offline”, “online” or you might see that he is viewing or replying to a specific question. You might as well see that he is just looking around in a specific section. \(\normalsize\color{blue}{\rm {You~~~can~~~ see~~~if~~~a~~~user~~~is~~~offline~~~by} }\) Clicking on the user's username, and below his picture you will see (in light blue) “offline”, “online” or you might see that he is viewing or replying to a specific question. You might as well see that he is just looking around in a specific section. \(\normalsize\color{purple }{\tt{ If~~~you~~~have~~~~any~~~questions~~~you}}\) \(\normalsize\color{purple}{\tt{can~~~message~~~me~~~or~~~anyone~~~else} }\) \(\normalsize\color{purple}{ \tt{ on~~~~this~~~~website.~~~~We~~~~will~~~be}}\) \(\normalsize\color{purple}{ \tt{ happy~~~to~~~respond~!~}}\)